if x=3 and x=-5, then form a quadratic equation. How can I do this to come out right? Thanks
Answer by arunpaul(104)

x=3 and x=-5,
when x =3
then x-3 =0 --- eq1
when x=-5
then x+5 =0 ----eq2
as the two equation gives 0 result then their multiplication must give 0 result
so (x-3) * (x+5) =0 (* sign indicate multiplication)
or x^2+5x-3x -15 =0 ( by expanding and multiplying the left hand side)
or x^2+2x-15 =0
this is the required answer (arun_662002@yahoo.co.in)